How to calculate the pressure, in atmospheres, of 2.40 moles helium gas in a 8.00 L container at 18 °C.
1
Expert's answer
2020-10-22T14:15:35-0400

pV=nRT,

where R is a gas constant (for pressure in atm we should take 0.082 L•atm/mol•K),

T = 18+273=291(K).

p=nRT/V

p=2.4×0.082×291/8=7.16(atm)